# Interpretable Concept-Based Memory Reasoning

### v1

@@ C.3.1 Concept interventions
To measure the effectiveness of concept interventions [4], we report task accuracy before and after replacing the concept predictions with their ground truth. In Table 4, we observe that CMR is responsive to concept interventions: After the interventions, CMR achieves perfect task accuracy, outperforming CEM and DCR. We leave a more extensive investigation of concept interventions for CMR to future work.
@@ C.3.3 Decoded rulebooks
For brevity, we drop the notation that shows whether the concept is related to the first or second digit. Firstly, consider the tasks $y_{0}$ and $y_{1}$. For these tasks, we are essentially missing all the concepts we need for the ground truth rules; we can only predict True by learning the rule that is the negation of all concepts (in the used rule notation, the empty rule), which CMR does. Then, to be able to predict False, it learns some arbitrary rules. Secondly, consider the tasks $y_{2}$ to $y_{10}$. As we are missing the concepts for digits 0 and 1, we cannot learn all ground truth rules for these tasks. For example, while CMR can (and does) still learn the rule $y_{4}\leftarrow 2\land 2$, it is impossible to learn a rule like $y_{4}\leftarrow 1\land 3$. Instead, CMR learns the rule $y_{4}\leftarrow 3$, which can be selected in case it needs to predict that $1+3=4$. Lastly, consider the tasks $y_{11}$ to $y_{18}$. CMR finds all ground truth rules, as they can be found. We show the learned rules in Figure 5.
@@
$y_{0}\leftarrow 4\land 7$
$y_{0}\leftarrow 4\land 8$
$y_{0}\leftarrow 8\land 9\land 9$
@@
$y_{1}\leftarrow 3\land 9$
$y_{1}\leftarrow 7$
$y_{1}\leftarrow 8$
@@
Table 9: Rulebook for CelebA ($n_{C}=1$).
@@
$(2)\quad\quad Black\_Hair\leftarrow\neg 5\_o\_Clock\_Shadow\land\neg Bags\_Under\_Eyes\land\neg Bald\land\neg Blurry\land\neg Bushy\_Eyebrows\land\neg Chubby$
$(3)\quad\quad Black\_Hair\leftarrow\neg Arched\_Eyebrows\land\neg Blurry$
@@
$(5)\quad\quad Male\leftarrow\neg 5\_o\_Clock\_Shadow\land\neg Bags\_Under\_Eyes\land\neg Bald\land\neg Blurry\land\neg Brown\_Hair\land\neg Chubby$
$(6)\quad\quad Male\leftarrow\neg Bald\land\neg Blurry$
@@
$(8)\quad\quad Wavy\_Hair\leftarrow\neg 5\_o\_Clock\_Shadow\land\neg Bags\_Under\_Eyes\land\neg Bald\land\neg Big\_Nose\land\neg Blurry\land\neg Bushy\_Eyebrows\land\neg Chubby$
$(9)\quad\quad Wavy\_Hair\leftarrow\neg Arched\_Eyebrows\land\neg Bald\land\neg Bangs\land\neg Blurry$
Table 11: Rulebook for CelebA ($n_{C}=37$) (not complete).
$(1)\quad\quad Black\_Hair\leftarrow 5\_o\_Clock\_Shadow\land Arched\_Eyebrows\land Bags\_Under\_Eyes\land Bald\land Bangs\land\neg Big\_Lips\land Big\_Nose\land Blond\_Hair\land Brown\_Hair\land Bushy\_Eyebrows\land Chubby\land\neg Double\_Chin\land Eyeglasses\land Goatee\land\neg Gray\_Hair\land Heavy\_Makeup\land\neg Mouth\_Slightly\_Open\land Mustache\land\neg Narrow\_Eyes\land\neg No\_Beard\land Oval\_Face\land Pale\_Skin\land\neg Pointy\_Nose\land Receding\_Hairline\land Rosy\_Cheeks\land Sideburns\land Straight\_Hair\land Wearing\_Hat\land Wearing\_Lipstick\land Wearing\_Necklace\land\neg Wearing\_Necktie\land Attractive\land\neg Young$
$(2)\quad\quad Black\_Hair\leftarrow 5\_o\_Clock\_Shadow\land Bags\_Under\_Eyes\land Bald\land\neg Big\_Lips\land Big\_Nose\land Blond\_Hair\land Brown\_Hair\land Bushy\_Eyebrows\land Chubby\land\neg Double\_Chin\land Eyeglasses\land Gray\_Hair\land Heavy\_Makeup\land\neg High\_Cheekbones\land\neg Mouth\_Slightly\_Open\land Mustache\land Narrow\_Eyes\land\neg No\_Beard\land\neg Oval\_Face\land Pale\_Skin\land\neg Pointy\_Nose\land Receding\_Hairline\land\neg Rosy\_Cheeks\land Straight\_Hair\land Wearing\_Lipstick\land\neg Wearing\_Necklace\land\neg Wearing\_Necktie\land\neg Young$
$(3)\quad\quad Black\_Hair\leftarrow\neg 5\_o\_Clock\_Shadow\land\neg Bald\land\neg Big\_Lips\land\neg Chubby\land\neg Double\_Chin\land\neg Eyeglasses\land\neg Goatee\land\neg Gray\_Hair\land\neg Mustache\land\neg Narrow\_Eyes\land\neg Pale\_Skin\land\neg Pointy\_Nose\land\neg Receding\_Hairline\land\neg Rosy\_Cheeks\land\neg Sideburns\land\neg Wearing\_Necklace\land\neg Wearing\_Necktie$
@@
$(9)\quad\quad Wavy\_Hair\leftarrow 5\_o\_Clock\_Shadow\land Arched\_Eyebrows\land Bags\_Under\_Eyes\land Bald\land Bangs\land\neg Big\_Lips\land Big\_Nose\land Blond\_Hair\land Blurry\land\neg Brown\_Hair\land\neg Bushy\_Eyebrows\land Chubby\land\neg Double\_Chin\land Eyeglasses\land\neg Gray\_Hair\land Heavy\_Makeup\land\neg Mouth\_Slightly\_Open\land\neg Mustache\land\neg Narrow\_Eyes\land\neg No\_Beard\land\neg Oval\_Face\land Pale\_Skin\land\neg Pointy\_Nose\land Receding\_Hairline\land Rosy\_Cheeks\land\neg Sideburns\land Straight\_Hair\land Wearing\_Lipstick\land\neg Wearing\_Necklace\land\neg Wearing\_Necktie\land Attractive\land\neg Young$
$(10)\quad\quad Wavy\_Hair\leftarrow\neg 5\_o\_Clock\_Shadow\land Arched\_Eyebrows\land Bags\_Under\_Eyes\land Bald\land Bangs\land Big\_Lips\land Big\_Nose\land Blond\_Hair\land Brown\_Hair\land Bushy\_Eyebrows\land Chubby\land\neg Double\_Chin\land Eyeglasses\land Goatee\land Heavy\_Makeup\land\neg No\_Beard\land\neg Oval\_Face\land\neg Pointy\_Nose\land Receding\_Hairline\land Rosy\_Cheeks\land\neg Sideburns\land\neg Smiling\land\neg Wearing\_Earrings\land\neg Wearing\_Necklace\land\neg Wearing\_Necktie\land\neg Young$
$(11)\quad\quad Wavy\_Hair\leftarrow\neg 5\_o\_Clock\_Shadow\land\neg Bags\_Under\_Eyes\land\neg Bald\land\neg Big\_Nose\land\neg Blurry\land\neg Bushy\_Eyebrows\land\neg Chubby\land\neg Double\_Chin\land\neg Eyeglasses\land\neg Goatee\land\neg Gray\_Hair\land\neg Mustache\land\neg Narrow\_Eyes\land\neg Pale\_Skin\land\neg Receding\_Hairline\land\neg Sideburns\land\neg Straight\_Hair\land\neg Wearing\_Hat\land\neg Wearing\_Necktie$
$(12)\quad\quad Wavy\_Hair\leftarrow\neg Bald\land\neg Mustache\land\neg Narrow\_Eyes\land\neg Pale\_Skin\land\neg Rosy\_Cheeks$
@@ Appendix D Code, licenses and resources
For our experiments, we implemented the models in Python 3.11.5 using open source libraries. This includes PyTorch v2.1.1 (BSD license) [45], PyTorch-Lightning v2.1.2 (Apache license 2.0), scikit-learn v1.3.0 (BSD license) [46] and xgboost v2.0.3 (Apache license 2.0). We used CUDA v12.4. Plots were made using Matplotlib v3.8.0 (BSD license) [47].

### v2

@@ C.3.1 Concept interventions
To measure the effectiveness of concept interventions [koh2020concept], we report task accuracy before and after replacing the concept predictions with their ground truth. In Table 4, we observe that CMR is responsive to concept interventions: After the interventions, CMR achieves perfect task accuracy, outperforming CEM and DCR. We leave a more extensive investigation of concept interventions for CMR to future work.
@@ C.3.3 Decoded rulebooks
For brevity, we drop the notation that shows whether the concept is related to the first or second digit. Firstly, consider the tasks $y_{0}$ and $y_{1}$. For these tasks, we are essentially missing all the concepts we need for the ground truth rules; we can only predict True by learning the rule that is the negation of all concepts (in the used rule notation, the empty rule), which CMR does. Then, to be able to predict False, it learns some arbitrary rules. Secondly, consider the tasks $y_{2}$ to $y_{10}$. As we are missing the concepts for digits 0 and 1, we cannot learn all ground truth rules for these tasks. For example, while CMR can (and does) still learn the rule $y_{4}\leftarrow 2\land 2$, it is impossible to learn a rule like $y_{4}\leftarrow 1\land 3$. Instead, CMR learns the rule $y_{4}\leftarrow 3$, which can be selected in case it needs to predict that $1+3=4$. Lastly, consider the tasks $y_{11}$ to $y_{18}$. CMR finds all ground truth rules, as they can be found. We show the learned rules in Figure LABEL:tab:rules_mnist_missing_seed_1.
@@
$y_{0}\leftarrow 6\land 7\land 8\land 9$
$y_{0}\leftarrow 7\land 4$
$y_{0}\leftarrow 8$
@@
$y_{1}\leftarrow 3$
$y_{1}\leftarrow 9\land 9$
@@
Table 9: Rulebook for CelebA ($n_{C}=1$) (seed 1).
@@
$(2)\quad\quad Black\_Hair\leftarrow\neg 5\_o\_Clock\_Shadow\land\neg Bags\_Under\_Eyes\land\neg Bald\land\neg Big\_Lips\land\neg Big\_Nose\land\neg Bushy\_Eyebrows\land\neg Chubby$
$(3)\quad\quad Black\_Hair\leftarrow\neg Arched\_Eyebrows\land\neg Blond\_Hair\land\neg Blurry$
@@
$(5)\quad\quad Male\leftarrow\neg 5\_o\_Clock\_Shadow\land\neg Bags\_Under\_Eyes\land\neg Bald\land\neg Blond\_Hair\land\neg Blurry\land\neg Brown\_Hair\land\neg Chubby$
$(6)\quad\quad Male\leftarrow\neg Bald\land\neg Blond\_Hair$
@@
$(8)\quad\quad Wavy\_Hair\leftarrow\neg 5\_o\_Clock\_Shadow\land\neg Bags\_Under\_Eyes\land\neg Bald\land\neg Big\_Nose\land\neg Bushy\_Eyebrows\land\neg Chubby$
$(9)\quad\quad Wavy\_Hair\leftarrow\neg Arched\_Eyebrows\land\neg Bald\land\neg Blond\_Hair\land\neg Blurry\land\neg Chubby$
Table 11: Rulebook for CelebA ($n_{C}=37$).
$(1)\quad\quad Black\_Hair\leftarrow 5\_o\_Clock\_Shadow\land Arched\_Eyebrows\land Bags\_Under\_Eyes\land Bald\land\neg Big\_Lips\land Chubby\land\neg Double\_Chin\land\neg Eyeglasses\land\neg Gray\_Hair\land Heavy\_Makeup\land\neg Mouth\_Slightly\_Open\land\neg No\_Beard\land Oval\_Face\land Pale\_Skin\land\neg Pointy\_Nose\land Receding\_Hairline\land\neg Rosy\_Cheeks\land\neg Wearing\_Necklace\land\neg Wearing\_Necktie$
$(2)\quad\quad Black\_Hair\leftarrow 5\_o\_Clock\_Shadow\land Arched\_Eyebrows\land Bags\_Under\_Eyes\land\neg Bald\land\neg Big\_Lips\land Big\_Nose\land Blond\_Hair\land\neg Bushy\_Eyebrows\land Chubby\land\neg Double\_Chin\land Eyeglasses\land Goatee\land\neg Gray\_Hair\land Heavy\_Makeup\land\neg Mouth\_Slightly\_Open\land Mustache\land\neg Narrow\_Eyes\land\neg No\_Beard\land Oval\_Face\land Pale\_Skin\land\neg Pointy\_Nose\land Receding\_Hairline\land Rosy\_Cheeks\land Straight\_Hair\land Wearing\_Hat\land Wearing\_Lipstick\land Wearing\_Necklace\land\neg Wearing\_Necktie\land Attractive\land\neg Young$
@@
$(4)\quad\quad Black\_Hair\leftarrow\neg Bald\land\neg Big\_Lips\land\neg Blurry\land\neg Chubby\land\neg Double\_Chin\land\neg Eyeglasses\land\neg Goatee\land\neg Gray\_Hair\land\neg Mustache\land\neg Narrow\_Eyes\land\neg Pale\_Skin\land\neg Pointy\_Nose\land\neg Receding\_Hairline\land\neg Rosy\_Cheeks\land\neg Sideburns\land\neg Wearing\_Necklace\land\neg Wearing\_Necktie$
$(5)\quad\quad Male\leftarrow 5\_o\_Clock\_Shadow\land Arched\_Eyebrows\land Bags\_Under\_Eyes\land\neg Bald\land Big\_Lips\land Big\_Nose\land Blond\_Hair\land Blurry\land Bushy\_Eyebrows\land Chubby\land\neg Double\_Chin\land Eyeglasses\land Goatee\land\neg Gray\_Hair\land Heavy\_Makeup\land High\_Cheekbones\land Mouth\_Slightly\_Open\land Mustache\land\neg Narrow\_Eyes\land\neg No\_Beard\land Oval\_Face\land Pale\_Skin\land\neg Pointy\_Nose\land Receding\_Hairline\land\neg Rosy\_Cheeks\land Straight\_Hair\land Wearing\_Hat\land Wearing\_Lipstick\land Wearing\_Necklace\land\neg Wearing\_Necktie\land Attractive\land\neg Young$
$(6)\quad\quad Male\leftarrow 5\_o\_Clock\_Shadow\land Arched\_Eyebrows\land\neg Bags\_Under\_Eyes\land\neg Big\_Lips\land Chubby\land\neg Double\_Chin\land\neg Eyeglasses\land\neg Goatee\land\neg Gray\_Hair\land High\_Cheekbones\land No\_Beard\land\neg Pointy\_Nose\land Receding\_Hairline\land\neg Rosy\_Cheeks\land\neg Sideburns\land\neg Wearing\_Necklace\land\neg Wearing\_Necktie$
$(7)\quad\quad Male\leftarrow\neg 5\_o\_Clock\_Shadow\land Arched\_Eyebrows\land\neg Bags\_Under\_Eyes\land\neg Bald\land Bangs\land Big\_Lips\land Big\_Nose\land Brown\_Hair\land Bushy\_Eyebrows\land\neg Chubby\land\neg Double\_Chin\land Eyeglasses\land Goatee\land Heavy\_Makeup\land Narrow\_Eyes\land\neg No\_Beard\land Oval\_Face\land Pale\_Skin\land Receding\_Hairline\land Wearing\_Earrings\land\neg Wearing\_Lipstick\land\neg Wearing\_Necklace\land\neg Wearing\_Necktie\land\neg Young$
$(8)\quad\quad Male\leftarrow\neg Bald\land\neg Blond\_Hair\land\neg Blurry\land\neg Gray\_Hair\land\neg Mustache\land\neg Narrow\_Eyes\land\neg Pale\_Skin\land\neg Receding\_Hairline\land\neg Rosy\_Cheeks\land\neg Wearing\_Necklace$
$(9)\quad\quad Wavy\_Hair\leftarrow 5\_o\_Clock\_Shadow\land Arched\_Eyebrows\land\neg Bags\_Under\_Eyes\land Bangs\land Big\_Lips\land Brown\_Hair\land Bushy\_Eyebrows\land\neg Chubby\land\neg Double\_Chin\land Eyeglasses\land Goatee\land Heavy\_Makeup\land\neg No\_Beard\land\neg Pointy\_Nose\land Receding\_Hairline\land Rosy\_Cheeks\land Sideburns\land\neg Wearing\_Earrings\land\neg Wearing\_Necklace\land\neg Wearing\_Necktie\land\neg Young$
$(10)\quad\quad Wavy\_Hair\leftarrow 5\_o\_Clock\_Shadow\land Arched\_Eyebrows\land\neg Big\_Lips\land Blond\_Hair\land Blurry\land\neg Brown\_Hair\land\neg Bushy\_Eyebrows\land Chubby\land\neg Double\_Chin\land Eyeglasses\land\neg Gray\_Hair\land Heavy\_Makeup\land\neg Mustache\land\neg Narrow\_Eyes\land\neg No\_Beard\land\neg Pointy\_Nose\land Receding\_Hairline\land Straight\_Hair\land Wearing\_Lipstick\land\neg Wearing\_Necklace\land\neg Wearing\_Necktie\land Attractive\land\neg Young$
$(11)\quad\quad Wavy\_Hair\leftarrow\neg 5\_o\_Clock\_Shadow\land\neg Bags\_Under\_Eyes\land\neg Bald\land\neg Big\_Nose\land\neg Blurry\land\neg Bushy\_Eyebrows\land\neg Chubby\land\neg Double\_Chin\land\neg Eyeglasses\land\neg Goatee\land\neg Gray\_Hair\land\neg Mustache\land\neg Narrow\_Eyes\land\neg Pale\_Skin\land\neg Receding\_Hairline\land\neg Sideburns\land\neg Straight\_Hair\land\neg Wearing\_Necklace\land\neg Wearing\_Necktie$
$(12)\quad\quad Wavy\_Hair\leftarrow\neg Bald\land\neg Chubby\land\neg Mustache\land\neg Narrow\_Eyes\land\neg Pale\_Skin\land\neg Receding\_Hairline\land\neg Rosy\_Cheeks\land\neg Wearing\_Necklace$
@@ Appendix D Code, licenses and resources
For our experiments, we implemented the models in Python 3.11.5 using open source libraries. This includes PyTorch v2.1.1 (BSD license) [paszke2019pytorch], PyTorch-Lightning v2.1.2 (Apache license 2.0), scikit-learn v1.3.0 (BSD license) [pedregosa2018scikitlearn] and xgboost v2.0.3 (Apache license 2.0). We used CUDA v12.4. Plots were made using Matplotlib v3.8.0 (BSD license) [Hunter:2007].